### v1

# Exact Quantum Capacity of Decohering Channels in Arbitrary Dimensions

Abstract
We derive an exact analytical expression for the quantum capacity of a broad class of decohering channels of the form $\Lambda(\rho)=(1-x)\rho+xD(\rho),$ where $D(\rho)$ denotes different forms of decoherence of the density matrix $\rho$. Unlike previous studies restricted to qubit systems or asymptotic bounds, our result holds for arbitrary finite-dimensional Hilbert spaces. We show that these channels are degradable for all $x$, leading to a closed-form, single-letter capacity formula that does not require regularization. Besides full decoherence, we extend our analysis to block-decohering channels, where decoherence occurs within orthogonal or overlapping subspaces. This family interpolates between identity channels and classical-quantum channels. Our findings clarify the role of coherence in quantum communication and provide rare exact benchmarks in the study of quantum channel capacities.
1 Introduction
Decoherence, the loss of quantum coherence due to environmental interactions, lies at the heart of quantum physics, governing the transition from quantum to classical behavior and imposing fundamental limits on quantum technologies [1, 2, 3]. It manifests as dephasing (phase randomization), amplitude damping (energy loss), or generalized noise, destroying superposition and entanglement both of which are essential for quantum computation , communication, and sensing[4, 5, 6]. While decoherence explains emergent classicality in macroscopic systems [7, 8, 2], its mitigation is critical for scalable quantum devices, where noise can destroy fragile quantum states. Understanding its precise impact on quantum information processing, particularly through the lens of channel capacities, remains a central challenge due to the superadditivity of coherent information [9, 10, 11, 12] and the need for regularization [13, 14, 15] in capacity calculations.
In this work, we address this challenge by deriving exact quantum capacities for a broad class of decohering channels, where noise is dominated by dephasing. While there are few channels for which quantum capacity is exactly known, like the bit flip channel [15] or the erasure channel [14], we identify a class of operationally important quantum channels, of the form
$\Lambda(\rho)=(1-x)\rho+xD(\rho),$
where $D(\rho)$ induces complete dephasing. Understandably, the capacity of these channels reduce to zero, when decoherence is complete at $x=1$. We then obtain the exact quantum capacity for family of channels of the form
$\Lambda_{k}(\rho)=(1-x)\rho+xD_{k}(\rho),$
where now block-decohrence [16, 17, 18, 19, 20, 21, 22, 23] occurs between orthogonal subspaces of dimension $k$ and is preserved in each subspace.
The surprize is when decoherence acts between subspaces rather than individual states: even under maximal noise ($x=1$), the channel retains non-zero capacity, if coherence persists within subspaces. This demonstrates that subspace-localized coherence, a form of "protected" quantum information, can defy complete erasure. Our explicit capacity formula quantifies this effect, and shows how decoherence-free subspaces [24] contribute in the capacity of this channel. These results establish rigorous benchmarks for fault-tolerant protocols and reveal hidden resources in partially coherent systems. Our results bridge foundational questions about the quantum-classical divide with practical constraints in quantum engineering, and offers tools to optimize information transmission in quantum networks and quantum memories.
To go beyond this decoherence-free regime, we derive the exact capacity for a third class of quantum channels of the form
$\Phi_{k}(\rho)=(1-x)\rho+xW_{k}(\rho),$
where $W_{k}$ destroys coherence between overlapping subspaces and there is no decoherence-free subspace. Even in this case, we see that capacity remains non-zero in the limit of strong decoherence.
Our method relies on an explict proof of degradability [25, 26] of all these channels, through a method which can set a method for constructing large families of degradable channels.
The structure of this paper is as follows: In section 2, we collect some known facts about the quantum capacity of quantum information and describe why it is do difficult to obtain exact results for capacities. In 3, we first study the case of fully decohering channel $\Lambda$, where the dimension of each subspace is $1$. After proving its degradability property we proceed to calculate its quantum capacity. We then go on in section 4 to consider the case where decoherence occurs between orthogonal subspaces of dimension $k$. The proofs and methods of calculation are similar to the case of $k=1$ with slight differences. Finally in section 5, we study the case of decoherence among overlapping subspaces. The basic resuts are shown in figures (1) and (2). The paper is concluded with a discussion.
2 Preliminaries on quantum capacities
A central concept in the study of quantum channels is the quantum capacity, which quantifies the maximum rate at which quantum information can be reliably transmitted through a channel. The quantum capacity is a key measure of channel performance [27, 28, 29], directly impacting the efficiency and reliability of quantum communication systems. However, calculating the quantum capacity is notoriously difficult, even for relatively simple channels like the depolarizing channel. This difficulty arises from the generic super-additivity property of most quantum channels. This is one of the most challenging aspects of quantum capacity. Unlike classical capacities, where the capacity of two independent channels is simply the sum of their individual capacities, the quantum capacity of the tensor product of two channels can be greater than the sum of the capacities of the individual channels. This behavior significantly complicates the calculation of quantum capacity because it requires considering not just single uses of the channel, but also the possibility of using entangled input states across multiple channel uses. As a result, the quantum capacity of a channel can only be determined by optimizing the coherent information, over an infinite number of uses of the channel [13], a task which in general is intractable. More precisely the quantum capacity of a channel ${\cal E}$ should be calculated via [29]:
$\begin{split}Q({\cal E})=\lim_{n\to\infty}\;\frac{1}{n}Q^{1}({\cal E}^{\otimes n}),\end{split}$ (1)
[…]
An important development in this direction was made by introducing the concepts of degradablity of quantum channels [25, 26].
Definition 1.
[25, 26] Consider a quantum channel ${\cal E}:A\to B$ and its complementary channel ${\cal E}^{c}:A\to E$. The channel ${\cal E}$ is said to be degradable if there exists a quantum channel $\mathcal{M}:B\to E$ such that $\mathcal{M}\circ{\cal E}(\rho)={\cal E}^{c}(\rho)$. Conversely it is called anti-degradable if there is a quantum channel $\mathcal{N}:B\to E$ such that $\mathcal{N}\circ{\cal E}^{c}(\rho)={\cal E}(\rho)$.
For a degradable channel the additivity property holds and we have $Q({\cal E})=Q^{1}({\cal E})$ [25], and in this case the calculation of the quantum capacity becomes a convex optimization problem. The capacity of an anti-degradable channel is zero.
What we will show in this letter is to prove that the decohering channels $\Lambda,\Lambda_{k}$ and $\Phi_{k}$ are all degradable for all values of $x$ in every dimension and use this property to find their capacity in closed form. The results are shown in figures (1) and (2).
3 The decohering channel
We consider the channel
$\Lambda(\rho)=(1-x)\rho+x\ D(\rho)$ (2)
where $x\in[0,1]$ and $D(\rho)=\sum_{i=1}^{d}\rho_{ii}|i\rangle\langle i|$ shows the decohered state of $\rho$. This channel shows the effect of decohering noise which happens with probability $x$. The first basic property of this channel, which will play a central role in proving the degradability of the channel is that
$D(\Lambda(\rho))=D(\rho).$ (3)
The Kraus operators of the channel are given by
$A_{0}=\sqrt{1-x}\ I\hskip 28.45274ptA_{i}=\sqrt{x}\ |i\rangle\langle i|.$ (4)
From the isometric extension, we find that the complementary channel of $\Lambda$ is given by
$\Lambda^{c}(\rho)=\sum_{i,j}Tr(A_{i}\rho A_{j}^{\dagger})\ |i\rangle\langle j|.$ (5)
Therefore we find from (4) and (5),
$\Lambda^{c}(\rho)=\begin{pmatrix}(1-x)Tr(\rho)&a\rho_{11}&a\rho_{22}&\cdot&a\rho_{dd}\\ a\rho_{11}&x\rho_{11}&0&0&0\\ a\rho_{22}&0&x\rho_{22}&0&0\\ \cdot&0&0&\cdot&0&\\ a\rho_{dd}&0&0&0&x\rho_{dd},\end{pmatrix}$ (6)
where $a=\sqrt{x(1-x)}$. The crucial property is that $\Lambda^{c}(Y)$ depends only on the diagonal entries of $Y$, that is
$\Lambda^{c}(Y)=\Lambda^{c}(D(Y)).$ (7)
This leads to the second basic property of the channel, namely
$\Lambda^{c}(\Lambda(\rho))=\Lambda^{c}\big{(}\ D(\Lambda(\rho))\ \big{)}.$ (8)
We now use combine the two basic properties of this channel (3) and (7) to prove that it is degradable. The reasoning is quite simple. We write
$\Lambda^{c}(\Lambda(\rho))=\Lambda^{c}(D(\Lambda(\rho)))=\Lambda^{c}(D(\rho)))=\Lambda^{c}(\rho),$ (9)
where in the first equality we have used (7), in the second, we have used (3) and then in the third equality, we have again used (7). Thus we have proved that the channel $\Lambda$ is degradable and $\Lambda^{c}$ in fact degrades $\Lambda$, that is
$\Lambda^{c}\circ\Lambda=\Lambda^{c}.$
Note that this is entirely different from self-complementary channels [39], i.e. those for which $\Lambda^{c}=\Lambda$. These channels are both degradable and anti-degradable and hence have zero capacity. The degradability of the decohering channel $\Lambda$, has now important consequences, namely that its full quantum capacity is the same as its one-shot capacity $Q^{(1)}(\Lambda).$ Moreover, degradability also implies that the coherent information $J(\Lambda,\rho)$ is a concave function of $\rho$. These two results greatly facilitates the exceedingly difficult task of optimization in eq. (1). First we can bypass the almost impossible task of regularization over infinitely many uses of the channel. Second we use concavity of the coherent information to do the optimization in a direct analytical way. To do this, we note that due to the Schur-Horn theorem in matrix theory, the diagonal entries of any Hermitian matrix are majorized by its eigenvalues. Thus we have
$d(A)\prec\lambda(A)$ (10)
which imediately leads to
$S(\Lambda(\rho))\leq S(D(\Lambda(\rho))).$ (11)
For the complementary channel we also know from $\Lambda^{c}\circ D=\Lambda^{c}$ that
$S(\Lambda^{c}(\rho))=S(\ \Lambda^{c}(D(\rho))\ ).$ (12)
These two results together imply that to maximize the coherent information, it is sufficient to only search the space of diagonal matrices $\rho_{d}$, namely
$J_{c}(\Lambda,\rho)\leq J_{c}(\Lambda,\rho_{d}).$ (13)
To find the state which maximizes the coherent information, we invoke the covariance property of the channel, namely
$\Lambda(X\rho X^{\dagger})=X\Lambda(\rho)X^{\dagger},$ (14)
where $\rho$ is any matrix (diagonal or not) and $X$ is the shift operator
$X=\sum_{i}|i+1\rangle\langle i|.$ (15)
Like any other channel [38], the covariance of the channel imposes the covariance of the complementary channel which in the present case takes the form
$\Lambda^{c}(X\rho X^{\dagger})=U_{x}\Lambda^{c}(\rho)U_{x}^{\dagger},$ (16)
where $U_{x}=|0\rangle\langle 0|\oplus X$. This means that $J_{c}(\rho)=J_{c}(X\rho X^{\dagger})$ for any density matrix. We now note that for diagonal density matrix $\rho_{d}$, it holds that
$\frac{1}{d}\sum_{i=1}^{d}X^{i}\rho_{d}{X^{\dagger}}^{i}=\frac{1}{d}I_{d}.$ (17)
Using this covariance and the concavity of the coherent information we can now write
$J_{c}(\frac{1}{d}I_{d})=J_{c}(\frac{1}{d}\sum_{i=1}^{d}X^{i}\rho_{d}{X^{\dagger}}^{i})\geq\frac{1}{d}\sum_{i=1}^{d}J_{c}(X^{i}\rho_{d}{X^{\dagger}}^{i})=\frac{1}{d}\sum_{i=1}^{d}J_{c}(\rho_{d})=J_{c}(\rho_{d})$ (18)
This means that the maximum is attained on the maximally mixed state and hence, due to unitality of the channel $\Lambda(\frac{I}{d})=\frac{I}{d}$, we have
$Q(\Lambda)=S(\frac{I}{d})-S(\Lambda^{c}(\frac{I}{d})).$ (19)
It is readily found that
$\Lambda^{c}(\frac{I}{d})=(1-x)|0\rangle\langle 0|+\frac{\sqrt{x(1-x)}}{d}\sum_{i=1}^{d}(|i\rangle\langle 0|+|0\rangle\langle i|)+\frac{x}{d}\sum_{i=1}^{d}|i\rangle\langle i|,$ (20)
or in explicit matrix form
$\Lambda^{c}(\frac{I}{d})=\begin{pmatrix}(1-x)&\xi\langle v_{0}|\\ \xi|v_{0}\rangle&\frac{x}{d}I_{d}\end{pmatrix}$ (21)
where $\xi=\sqrt{\frac{x(1-x)}{d}}$, $|v_{0}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}|i\rangle$ is the maximally coherent state and $I_{d}$ is the $d-$ dimensional identity matrix. The eigenvalues of this matrix are
$\{0,\frac{d+x-dx}{d},\frac{x}{d}\}$
where the last eigenvalue has degeneracy of $d-2$, (i.e. a total of $d$ eigenvalues). Collecting all terms, we find the quantum capacity of the channel $\Lambda_{x}$ in closed form as
$Q(\Lambda_{x})=\log d+\frac{d-1}{d}x\log\frac{x}{d}+(1-x+\frac{x}{d})\log(1-x+\frac{x}{d}).$ (22)
This is the exact capacity of the channel $\Lambda$ which smoothly drops from $\log_{2}d$ for the identity channel to 00 for the fully decohering channel, as expected. When $x=1$, the channel output is a classical mixture and has no superposition.
4 Quantum Capacity of Block-decohering channels
We now extend this method to the channel in which decoherence occurs between subspaces of dimension $k$ and coherence in each subspace is maintained. Such a channel is defined by the map
$\Lambda_{k}(\rho)=(1-x)\rho+xD_{k}(\rho).$ (23)
Here $D_{k}$ depicts the decoherence map between blocks
$D_{k}(\rho)=\begin{pmatrix}\rho_{1}&\cdot&\cdot&\cdot&\cdot\\ \cdot&\rho_{2}&\cdot&\cdot&\cdot\\ \cdot&\cdot&\rho_{3}&\cdot&\cdot\\ \cdot&\cdot&\ &\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\rho_{r}\end{pmatrix}$ (24)
where each $\rho_{i}$ is the projector of the density matrix onto the $i-th$ subspace of dimension $k$. The dimension of the full Hilbert space is taken to be $d=kr$, where $k$ is the dimension of each subspace and $r$ is the number of orthogonal subspaces, i.e. number of blocks. As a map $D_{k}(\rho)$ is given by
$D_{k}(\rho)=\sum_{i}\pi_{i=1}^{r}\rho\pi_{i},$ (25)
where $\pi_{1}=|1\rangle\langle 1|+\cdots|k\rangle\langle k|$, $\pi_{2}=|k+1\rangle\langle k+1|+\cdots|2k\rangle\langle 2k|$ and so on. The complementary channel now becomes
$\Lambda_{k}^{c}(\rho)=\begin{pmatrix}(1-x)Tr(\rho)&aTr(\rho_{1})&aTr(\rho_{2})&\cdot&aTr(\rho_{r})\\ aTr(\rho_{1})&xTr(\rho_{1})&0&0&0\\ aTr(\rho_{2})&0&xTr(\rho_{2})&0&0\\ \cdot&0&0&\cdot&0&\\ aTr(\rho_{r})&0&0&0&xTr(\rho_{r}),\end{pmatrix}$ (26)
where $a=\sqrt{x(1-x)}$. Both of the basic properties (3) and (7) hold also for this channel from which we find that this channel is also degradable. Moreover, the Schur-Horn theorem is used to show that optimization can be performed by searching only among the diagonal density matrices $\rho_{d}$. The channel now has an extra symmetry, that is it is covariance under two types of unitary operations, that is:
$\Lambda_{k}(\mathbb{X}\rho\mathbb{X})=\mathbb{X}\Lambda_{k}(\rho)\mathbb{X},\hskip 28.45274pt\Lambda_{k}(\Omega\rho\Omega)=\Omega\Lambda_{k}(\rho)\Omega,$ (27)
where
$\mathbb{X}=\begin{pmatrix}X_{k}&\cdot&\cdot&\cdot&\cdot\\ \cdot&X_{k}&\cdot&\cdot&\cdot\\ \cdot&\cdot&X_{k}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&X_{k}\end{pmatrix}$ (28)
which affects shift in each subspace and
$\Omega=\begin{pmatrix}\cdot&\cdot&\cdot&\cdot&I_{k}\\ \ I_{k}&\cdot&\cdot&\cdot&\cdot\\ \cdot&I_{k}&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&I_{k}&\cdot\end{pmatrix}$ (29)
which affects shifts among subspaces. We now use this symmetry to show that for an arbitrary diagonal matrix,
$\frac{1}{d}\sum_{i=1}^{k}\sum_{j=1}^{r}U_{ij}\rho U_{ij}^{\dagger}=\frac{I}{d},$ (30)
where $U_{ij}=\mathbb{X}^{i}\ \Omega^{j}$. In the same way as for the channel $\Lambda_{1}$, concavity of the coherent information now yields that
$Q(\Lambda_{k})=S(\Lambda_{k}(\frac{I}{d}))-S(\Lambda_{k}^{c}(\frac{I}{d}))$ (31)
where $d=rk$ is the dimension and $\Lambda_{k}^{c}$ has exactly the same form as (22), with $d$ replaced by $r$. The final result is given by
$Q(\Lambda_{k})=\log k+\log r+\frac{r-1}{r}x\log\frac{x}{r}+(1-x+\frac{x}{r})\log(1-x+\frac{x}{r}).$ (32)
The capacity of the channels at $x=0$ is given by $Q(\Lambda(x=0))=\log_{2}d$ and drops to $Q(\Lambda(x=1))=\log_{2}k$ This is a reflection of the fact that we can encode the input states into those subspaces which are not decohered. Plots of capacities of these channels is shown in figure (1) for the case where dimension $d=12$ and the number of blocks increase from $k=1$ (complete coherence) to $k=12$ (no coherence).
Figure 1: The quantum capacities of the block-decohering channels $\Lambda_{k}$ for $d=12$ for various values of $k$. For each $k$, the number of blocks is given by $r=\frac{12}{k}$. It is seen that even when $x=1$, full decoherence among subspaces, allows a nonzero quantum capacity. This is the capacity of quantum information transfer,when a suitable encoding into decoherence-free subspaces is used.
5 Quantum Capacity of Weakly-Decohering Channels
The non-zero capacity of block-decohering channels in the previous section may be ascribed to the presence of decoherence-free subspaces, namely the $k-$ dimensional decoherence-free subspaces which are not affected by the decohering noise. As we will see in this section, in the absence of these decoherence-free subspaces, and provided that some degree of coherence exists, the channel can still have non-zero quantum capacity. In this way, we are considering a noise which is not so strong to anihillate all the off-diagonal elements. Rather, we assume an equal number of minor diagonals above and below the main diagonal are also kept intact while all the other off-diagonal elements vanish. As an example, for $d=6$, an example of such weak decoherence is the following:
$\rho\longrightarrow W_{2}(\rho):=\begin{pmatrix}\rho_{11}&\rho_{12}&0&0&0&\rho_{16}\\ \rho_{21}&\rho_{22}&\rho_{23}&0&0&0\\ 0&\rho_{32}&\rho_{33}&\rho_{34}&0&0\\ 0&0&\rho_{43}&\rho_{44}&\rho_{45}&0\\ 0&0&0&\rho_{54}&\rho_{55}&\rho_{56}\\ \rho_{61}&0&0&0&\rho_{65}&\rho_{66}\end{pmatrix}.$
There is a related concept of coherence, called $k-$ coherence [22, 19], which is somewhat more general than this, but we will not consider it here. The original notion of incoherence [40], which defines incoherent states as diagonal density matrices in a specific basis, has been generalized to multi-level or k-coherence[19]. In this generalized setting, a state $|\psi\rangle=\sum_{i=1}^{d}\psi_{i}|i\rangle$ is said to have coherence at level k, if exactly $k$ of the coefficients $\psi_{i}$ are non-zero. Thus an incoherent state has coherence at level $1$, and a state like $|\psi\rangle=a|0\rangle+b|1\rangle$ in a $d-$ dimensional Hilbert space has coherence at level $2$ and so on. The generalization to mixed states is done by defining the states with coherence level $k$ to be the convex combination of all pure states whose coherence level is less than or equal to $k$. The notion of weak-coherence that we adopt here is different in that it assigns an order to the basis states. For example in the example (), If for some physical reason, say the photon number in each state, we order the states as $|0\rangle,|1\rangle,|2\rangle,\cdots$, two states of the form $|0\rangle+|1\rangle$ and $|1\rangle+|2\rangle$ are equally affected by decoherence, but a state like $|1\rangle+|3\rangle$ completely loses its coherence.
In this section, we consider decohering channels which with a certain probability decohere the input state to level-$k$ coherent states. Thats, we consider channesl of the form
$\Phi_{k}(\rho)=(1-x)\rho+xW_{k}(\rho),$ (33)
where $W_{k}(\rho)$ is a weakly coherent density matrix at level $k$. Physically this means that the channel mitigates all coherences among the basis states which are further apart than $k$ in an ordered set of indices $\{|1\rangle,|2\rangle,\cdots|d\rangle\}$. The Kraus operators of this channel are:
$A_{0}=\sqrt{1-x}I,\hskip 28.45274ptA_{i}=\sqrt{\frac{x}{k}}\sum_{l=0}^{k-1}|i+l\rangle\langle i+l|.$ (34)
In other words, each $A_{i}$ is proportional to the projection operator on a $k-$ dimensional subspace, but this time, in contrast to the previous section, the subspaces are not orthogonal and they overlap on smaller subspaces. We first note that the basic property (3) still holds here, that is
$D(\Phi_{k}(\rho))=\Phi_{k}(D(\rho))=D(\rho).$ (35)
For the second property (7), we show that for the channel $\Phi_{k}$, the complementary channel $\Phi_{k}^{c}$ still depends on the diagonal elements of $\rho$. To see this, it is sufficient to note that the Kraus operators $I$, $A_{i}$ and the products $A_{i}A_{j}$ are all projectors on subspaces and hence extarct the diagonal elements of $\rho$, when inserted in equation (5) for deriving the complementary channel. Therefore we have
$\Phi_{k}^{c}(\rho)=\Phi_{k}^{c}(D(\rho)).$ (36)
Combining these two equations, as before, we find
$\Phi_{k}^{c}(\Phi_{k}(\rho))=\Phi_{k}^{c}(D(\Phi_{k}(\rho)))=\Phi_{k}^{c}(D(\rho))=\Phi_{k}^{c}(\rho),$ (37)
which shows that the channel $\Phi_{k}$ is also degradable. Invoking Schur-Horn theorem as before, we need only search among the diagonal density matrices and noting the symmetry of the channel
$\Phi_{k}(X\rho X^{\dagger})=X\Phi_{k}(\rho)X^{\dagger},$ (38)
where $X$ is the shift operator on the whole space, i.e. $X=\sum_{i=1}^{d}|i+1\rangle\langle i|$, we find from the concave property of the coherent information that the maximum value is obtained at $\rho_{d}=\frac{I}{d}$. Thus we have
$Q_{k}(\Phi)=I_{c}(\Phi_{k},\frac{I}{d}).$
We first calculate the following elements:
[…]
where due to the period condition of indices $s\equiv s+d$, we find after a close examination that $b_{s}$ is given by
$b(s)=\begin{cases}k-s&\text{if }\ \ 0\leq s\leq k-1,\\ 0&\text{if }\ \ k\leq s\leq d-k,\\ s+k-d&\text{if}\ \ d-k+1\leq s\leq d-1.\end{cases}$ (42)
Thereofore we find that
$\left[\begin{array}[]{c|c}1-x&\xi\sqrt{k}\bra{v_{0}}\\ \hline\cr\xi\sqrt{k}\ket{v_{0}}&xB\end{array}\right]$ (43)
where again $\xi=\sqrt{\frac{x(1-x)}{d}}$ and |$\ket{v_{0}}=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\ket{i}$. and the matrix $B$ is a ciruclant matrix given by
$B=\mathrm{Circ}(b_{0},b_{1},\dots,b_{d-1})=\begin{bmatrix}b_{0}&b_{1}&b_{2}&\cdots&b_{d-2}&b_{d-1}\\ \\ b_{d-1}&b_{0}&b_{1}&\cdots&b_{d-3}&b_{d-2}\\ \\ b_{d-2}&b_{d-1}&b_{0}&\cdots&b_{d-4}&b_{d-3}\\ \\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ \\ b_{1}&b_{2}&b_{3}&\cdots&b_{d-1}&b_{0}\end{bmatrix}.$ (44)
The matrix $B$, being circulant, has the structure
$B=\sum_{s=0}^{d-1}b_{s}X^{s},$
where $X$ is the shift operator. The eigenvalues of $X$ are roots of unity in the form $\omega^{m}$ corresponding to the eigenvector $|v_{m}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\omega^{-im}|i\rangle$, ($m=0,1,\cdots d-1$), where $\omega=e^{\frac{2\pi i}{d}}$. Thus the eigenvalues of $B$ are of the form
$\lambda_{m}=\sum_{s=0}^{d-1}b_{s}\omega^{ms},\hskip 28.45274ptm=0,1,2,\cdots d-1.$ (45)
In view of (42), this turns out to be
$\lambda_{m}=\frac{1}{dk}\left[\frac{\sin{\left(\frac{k\pi m}{d}\right)}}{\sin{\left(\frac{\pi m}{d}\right)}}\right]^{2}.$ (46)
To find the quantum capacity of $\Phi_{k}$, we now need to diagonalize the matrix ${\cal M}:=\Phi_{k}^{c}(\frac{I}{d})$, as given in (43). We note that $|v_{0}\rangle$ is an eigenvector of $X$ and hence of $B$ with eigenvalue equal to $\lambda_{0}=\frac{k}{d}$. To find the spectrum of this matrix, we note that for any eigenvector $|v_{m\neq 0}\rangle$, the vector
$\begin{bmatrix}0\\ |v_{m}\rangle\end{bmatrix}$ (47)
is an eigenvector of ${\cal M}$ with eigenvalue equal to $x\lambda_{m}$. The other two eigenvectors lie in the span of |$\ket{v_{0}}$ and |$\ket{0}$:
${\cal M}(\alpha\ket{0}+\ket{v_{0}})=(\alpha(1-x)+d\zeta)\ket{0}+(\alpha\zeta+\frac{k}{d}x)\ket{v_{0}}$ (48)
The corresponding eigenvalue equation becomes, the solution of which yields two eigenvalues, namely
$\{0,\ \ \frac{kx}{d}+(1-x)\}.$
As a result, the set of eigenvalues of ${\cal M}=\Phi^{c}(\frac{I}{d})$ is:
$\{x\lambda_{1},x\lambda_{2},\ldots,x\lambda_{d-1}\}\cup\left\{0,\frac{kx}{d}+(1-x)\right\}$
where $\lambda_{m}$ is given in (46). This leads to the final expression of the quantum capacity, namely We can now evaluate the quantum capacity of $\Phi$:
$Q(\Phi_{k})=I_{c}\left(\Phi,\frac{I}{d}\right)=\log{d}+\sum_{i=1}^{d-1}x\lambda_{i}\log(x\lambda_{i})+\left[\frac{kx}{d}+(1-x)\right]\log{\left[\frac{kx}{d}+(1-x)\right]}$ )
Figure 2: Coherent information $I_{c}\bigl{(}\Phi,\tfrac{I}{d}\bigr{)}$ versus $x$ for $d=12$ and $k\in\{1,2,3,4,6\}$.
As special cases we note that when $k=1$, this capacity reduces to that of (22), as expected. Also for $x=0$, or when $k=d$, for which the channel is the identity channel, the capacity is given by $Q(\Phi_{k})=\log_{2}d.$ Note that for the strongest level of noise $(x=1)$, the quantum capacity is no longer given by $\log_{2}k$. The capacities of these channels for $d=12$ and for various degrees of decoherence, determined by $k$ are ploted in figure (2). It is worth nothing that even at full decoherence, the channel has non-zero capacity, despite having no decoherence-free subspace.
6 Discussion
We have derived the exact quantum capacity for a general class of decohering channels $\Lambda_{k}(\rho)=(1-x)\rho+xD_{k}(\rho)$. While exact quantum capacities are known for certain qubit Pauli channels (e.g., bit-flip with two Kraus operators), their higher-dimensional generalizations typically lack closed-form solutions due to superadditivity. In contrast, our decohering channels $\Lambda_{k}$, though structurally distinct from Pauli models. admit exact closed-form expressions for capacities for all values of $k$, $d$ and $x$.
Our results show that that even under maximal decoherence, ($x=1$) the channel retains non-zero capacity when coherence is preserved within subspaces. The capacity of course depends on the dimension of these subspaces. This challenges the intuition that complete dephasing destroys all quantum information, showing instead that structured noise can leave "islands" of coherence intact. Therefore, survival of quantum capacity under maximal decoherence (when $x=1$), suggests that one can protect information by encoding it into appropriate decoherence-free [24] subspaces.
The capacity formula explicitly quantifies how increasing the subspace dimension $k$ interpolates between classical and quantum transmission limits, with $k=1$ (full dephasing) and $k=d$ (th identity channel) as extremal cases. This dimensional hierarchy provides a practical toolkit for optimizing fault-tolerant encodings: i.e. when the type of coherence is known, one can use suitable encoding to protect quantum information. Our work thus establishes rigorous benchmarks for quantum error correction and underscores the role of subspace-localized coherence as a resource in noisy quantum systems.
Our work opens up several new directions in the problem of capacity of quantum channels. The first one is to use these results for obtaining upper or lower bounds for more general models of noise in arbitrary dimensions, i.e. when incomplete damping occurs in off-diagonal terms. The second avenue is based on the observation that the two crucial properties of this channel, namely $\Lambda\circ D=D$ and $\Lambda^{c}\circ D=\Lambda^{c}$, which have led to degradability of the channel $\Lambda$. One can ask whether there is a systematic way to find other maps ${\cal E}$, and ${\cal N}$, such that ${\cal E}\circ{\cal N}={\cal N}$ and ${\cal E}^{c}\circ{\cal D}={\cal E}^{c}$. One can see if these channels can activate superactivation [41, 42, 43] with zero-capacity channels. Finally The exact solutions we provide open avenues for designing noise-resilient quantum memories and networks, bridging fundamental limits with engineering constraints in the NISQ era [44, 45, 46, 47].
Acknowledgement: V.K. Thanks Dariusz Chruscinski, Saverio Pascazio, Gniewko Sarbicki and Mario Ziman for their valuable comments, when this work was presented in the 56th Symposium in Mathematical Physics in Torun, Poland.