### v2

# Exact Quantum Capacity of Decohering Channels in Arbitrary Dimensions

Abstract
We derive exact analytical expressions for the quantum capacity of a broad class of decohering quantum channels of the form $\Lambda(\rho)=(1-x)\rho+xD(\rho)$, where $D(\rho)$ represents a structured decoherence process. These channels are shown to be degradable for all noise parameters and in arbitrary dimensions, yielding closed-form, single-letter capacity formulas. Our analysis includes fully decohering, block-decohering, and weakly decohering channels, the latter involving coherence preservation within overlapping subspaces. Surprisingly, even under maximal decoherence, the channel may retain nonzero capacity due to residual coherence structure. These results provide quantitative role for decoherence-free and partially coherent subspaces in preserving quantum information, offering guidance for encoding strategies in quantum memories and fault-tolerant quantum communication systems.
1 Introduction
The capacity of a quantum channel to transmit information is a central figure of merit in quantum communication and quantum information processing [1, 2, 3]. Yet, despite its conceptual clarity, calculating the quantum capacity remains notoriously difficult, even for seemingly simple channels. This difficulty stems from the need to optimize the coherent information over multiple uses of the channel and the lack of additivity in general [4, 5, 6], rendering most capacities intractable. Exact capacity formulas are known only for a handful of special cases, such as the erasure channel [7], amplitude damping in limited regimes in low dimensions [8, 9, 10], and a few qubit Pauli channels [11]. In higher-dimensional systems, exact results are almost nonexistent [12, 13]. Even lower and upper bounds can be difficult [8, 14, 15, 16, 17, 18, 19, 20] to compute due to nonconvexities in the coherent information. Thus, any exact result in this domain not only deepens our theoretical understanding but also offers valuable benchmarks for real-world implementations. In this Letter, we present a broad family of quantum channels for which the quantum capacity can be computed exactly and in closed form. These channels take the form $\Lambda(\rho)=(1-x)\rho+xD(\rho)$, where $D(\rho)$ models structured decoherence [21, 22, 23, 24, 25]. Remarkably, we prove that these channels are degradable in all dimensions and for all values of the noise parameter x. This includes fully decohering channels, block-decohering models that preserve coherence within subspaces, and weakly decohering channels where coherence survives across overlapping modes. We show that even under strong noise $(x=1)$, these channels may retain nonzero quantum capacity—provided certain coherence structures are preserved [26, 27, 28, 29, 30, 25]. This finding challenges the common intuition that strong decoherence always implies complete quantum information loss.
We first consider the operationally important quantum channels, of the form
$\Lambda(\rho)=(1-x)\rho+xD(\rho),$
where $D(\rho)$ induces complete dephasing, then we study channels of the form
$\Lambda_{k}(\rho)=(1-x)\rho+xD_{k}(\rho),$
where now block-decoherence [31, 32, 33, 25, 34, 35, 36, 37] occurs between orthogonal subspaces of dimension $k$ and coherence is preserved in each subspace. The surprise is when decoherence acts between subspaces rather than individual states: even under maximal noise ($x=1$), the channel retains non-zero capacity, if coherence persists within subspaces. This demonstrates that subspace-localized coherence, a form of "protected" quantum information, can defy complete erasure. Our explicit capacity formula quantifies this effect, and shows how decoherence-free subspaces [38] contribute in the capacity of this channel. To go beyond these types of direct sum channels, we proceed to derive the exact capacity for a third class of quantum channels of the form
$\Phi_{k}(\rho)=(1-x)\rho+xW_{k}(\rho),$
where $W_{k}$ destroys coherence between overlapping subspaces in such a way that there is no decoherence-free subspace. Even in this case, we see that capacity remains non-zero in the limit of strong decoherence.
Our method relies on an explicit proof of degradability [8, 39] of all these channels, through a method which can set a new method for constructing large families of degradable channels.
The structure of this paper is as follows: In section 2, we collect some known facts about the quantum capacity of quantum information and describe why it is so difficult to obtain exact results for capacities. In 3, we first study the case of fully decohering channel $\Lambda$, where the dimension of each subspace is $1$. After proving its degradability property we proceed to calculate its quantum capacity. We then go on in section 4 to consider the case where decoherence occurs between orthogonal subspaces of dimension $k$. The proofs and methods of calculation are similar to the case of $k=1$ with slight differences. Finally in section 5, we study the case of decoherence among overlapping subspaces. The basic results are shown in figures (1) and (2). The paper is concluded with a discussion.
2 Preliminaries on quantum capacities
A central concept in the study of quantum channels is the quantum capacity, which quantifies the maximum rate at which quantum information can be reliably transmitted through a channel. The quantum capacity is a key measure of channel performance [40, 41, 42], directly impacting the efficiency and reliability of quantum communication systems. However, calculating the quantum capacity is notoriously difficult, even for relatively simple channels like the depolarizing channel. This difficulty arises from the generic super-additivity property of most quantum channels. This is one of the most challenging aspects of quantum capacity. Unlike classical capacities, where the capacity of two independent channels is simply the sum of their individual capacities, the quantum capacity of the tensor product of two channels can be greater than the sum of the capacities of the individual channels. This behavior significantly complicates the calculation of quantum capacity because it requires considering not just single uses of the channel, but also the possibility of using entangled input states across multiple channel uses. As a result, the quantum capacity of a channel can only be determined by optimizing the coherent information, over an infinite number of uses of the channel [43], a task which in general is intractable. More precisely the quantum capacity of a channel ${\cal E}$ should be calculated via [42]:
$\begin{split}Q({\cal E})=\lim_{n\to\infty}\;\frac{1}{n}Q^{1}({\cal E}^{\otimes n}),\end{split}$ (1)
[…]
An important development in this direction was made by introducing the concepts of degradability of quantum channels [8, 39].
Definition 1.
[8, 39] Consider a quantum channel ${\cal E}:A\to B$ and its complementary channel ${\cal E}^{c}:A\to E$, where $A$ and $B$ denote the domain and target systems and $E$ the auxiliary system playing the role of environment. The channel ${\cal E}$ is said to be degradable if there exists a quantum channel $\mathcal{M}:B\to E$ such that $\mathcal{M}\circ{\cal E}={\cal E}^{c}$. Conversely it is called anti-degradable if there is a quantum channel $\mathcal{N}:E\to B$ such that $\mathcal{N}\circ{\cal E}^{c}={\cal E}$.
For a degradable channel the additivity property holds and we have $Q({\cal E})=Q^{1}({\cal E})$ [8], and in this case the calculation of the quantum capacity becomes a convex optimization problem. The capacity of an anti-degradable channel is zero.
What we will show in this paper is to prove that the decohering channels $\Lambda,\Lambda_{k}$ and $\Phi_{k}$ are all degradable for all values of $x$ in every dimension and use this property to find their capacities in closed form. The results are shown in figures (1) and (2).
3 The fully-decohering channel $\Lambda$
We consider the channel
$\Lambda(\rho)=(1-x)\rho+x\ D(\rho)$ (2)
where $x\in[0,1]$ and $D(\rho)=\sum_{i=1}^{d}\rho_{ii}|i\rangle\langle i|$ shows the fully decohered state of $\rho$. This channel shows the effect of decohering noise which happens with probability $x$. The first basic property of this channel, which will play a central role in proving the degradability of the channel is that for all input states $\rho$, $D(\Lambda(\rho))=D(\rho).$ or simply
$D\circ\Lambda=D.$ (3)
The Kraus operators of the channel are given by
$A_{0}=\sqrt{1-x}\ I\hskip 28.45274ptA_{i}=\sqrt{x}\ |i\rangle\langle i|.$ (4)
From the isometric extension, we find that the complementary channel of $\Lambda$ is given by
$\Lambda^{c}(\rho)=\sum_{i,j}Tr(A_{i}\rho A_{j}^{\dagger})\ |i\rangle\langle j|.$ (5)
Therefore we find from (4) and (5),
$\Lambda^{c}(\rho)=\begin{pmatrix}(1-x)Tr(\rho)&a\rho_{11}&a\rho_{22}&\cdot&a\rho_{dd}\\ a\rho_{11}&x\rho_{11}&0&0&0\\ a\rho_{22}&0&x\rho_{22}&0&0\\ \cdot&0&0&\cdot&0&\\ a\rho_{dd}&0&0&0&x\rho_{dd},\end{pmatrix}$ (6)
where $a=\sqrt{x(1-x)}$. The second crucial property is that $\Lambda^{c}(Y)$ depends only on the diagonal entries of $Y$, that is, for any input state $\rho$
$\Lambda^{c}=\Lambda^{c}\circ D.$ (7)
The combination of (3) and (7)leads to the following chain of identities
$\Lambda^{c}\circ\Lambda=(\Lambda^{c}\circ D)\circ\Lambda=\Lambda^{c}\circ(D\circ\Lambda)=\Lambda^{c}\circ D=\Lambda^{c},$ (8)
where in the first equality we have used (7), in the second, we have used (3) and then in the third equality, we have again used (7). Thus we have proved that the channel $\Lambda$ is degradable and $\Lambda^{c}$ in fact degrades $\Lambda$, that is
$\Lambda^{c}\circ\Lambda=\Lambda^{c}.$
Note that this is entirely different from self-complementary channels [57], i.e. those for which $\Lambda^{c}=\Lambda$. These channels are both degradable and anti-degradable and hence have zero capacity. The degradability of the decohering channel $\Lambda$, has now important consequences, namely that its full quantum capacity is the same as its one-shot capacity $Q^{(1)}(\Lambda).$ Moreover, degradability also implies that the coherent information $J(\Lambda,\rho)$ is a concave function of $\rho$. These two results greatly facilitate the exceedingly difficult task of optimization in eq. (1). First we can bypass the almost impossible task of regularization over infinitely many uses of the channel. Second we use concavity of the coherent information to do the optimization in a direct analytical way. To do this, we note that due to the Schur-Horn theorem in matrix theory, the diagonal entries of any Hermitian matrix are majorized by its eigenvalues. Thus we have
$d(A)\prec\lambda(A)$ (9)
which immediately leads to
$S(\Lambda(\rho))\leq S(D(\Lambda(\rho))).$ (10)
For the complementary channel we also know from $\Lambda^{c}\circ D=\Lambda^{c}$ that
$S(\Lambda^{c}(\rho))=S(\ \Lambda^{c}(D(\rho))\ ).$ (11)
These two results together imply that to maximize the coherent information, it is sufficient to only search the space of diagonal matrices $\rho_{d}$, namely
$I_{c}(\Lambda,\rho)\leq I_{c}(\Lambda,\rho_{d}).$ (12)
[…]
$\Lambda(X\rho X^{\dagger})=X\Lambda(\rho)X^{\dagger},$ (13)
where $\rho$ is any matrix (diagonal or not) and $X$ is the shift operator
$X=\sum_{i}|i+1\rangle\langle i|.$ (14)
Like any other channel [55], the covariance of the channel imposes the covariance of the complementary channel which in the present case takes the form
$\Lambda^{c}(X\rho X^{\dagger})=U_{x}\Lambda^{c}(\rho)U_{x}^{\dagger},$ (15)
where $U_{x}=|0\rangle\langle 0|\oplus X$. This means that $I_{c}(\Lambda,\rho)=I_{c}(\Lambda,X\rho X^{\dagger})$ for any density matrix. We now note that for diagonal density matrix $\rho_{d}$, it holds that
$\frac{1}{d}\sum_{i=1}^{d}X^{i}\rho_{d}{X^{\dagger}}^{i}=\frac{1}{d}I_{d}.$ (16)
Using this covariance and the concavity of the coherent information we can now write
$I_{c}(\Lambda,\rho_{d})=\frac{1}{d}\sum_{i=1}^{d}I_{c}(\Lambda,\rho_{d})=\frac{1}{d}\sum_{i=1}^{d}I_{c}(\Lambda,X^{i}\rho_{d}{X^{\dagger}}^{i})\leq I_{c}(\Lambda,\frac{1}{d}\sum_{i=1}^{d}X^{i}\rho_{d}{X^{i}}^{\dagger})=I_{c}(\Lambda,\frac{1}{d}I_{d})$ (17)
This means that the maximum is attained on the maximally mixed state and hence, due to unitality of the channel $\Lambda(\frac{I}{d})=\frac{I}{d}$, we have
$Q(\Lambda)=S(\frac{I}{d})-S(\Lambda^{c}(\frac{I}{d})).$ (18)
It is readily found that
$\Lambda^{c}(\frac{I}{d})=(1-x)|0\rangle\langle 0|+\frac{\sqrt{x(1-x)}}{d}\sum_{i=1}^{d}(|i\rangle\langle 0|+|0\rangle\langle i|)+\frac{x}{d}\sum_{i=1}^{d}|i\rangle\langle i|,$ (19)
or in explicit matrix form
$\Lambda^{c}(\frac{I}{d})=\begin{pmatrix}(1-x)&\xi\langle v_{0}|\\ \xi|v_{0}\rangle&\frac{x}{d}I_{d}\end{pmatrix}$ (20)
where $\xi=\sqrt{\frac{x(1-x)}{d}}$, $|v_{0}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}|i\rangle$ is the maximally coherent state and $I_{d}$ is the $d-$ dimensional identity matrix. The eigenvalues of this matrix are
$\{0,\frac{d+x-dx}{d},\frac{x}{d}\}$
where the last eigenvalue has degeneracy of $d-2$, (i.e. a total of $d$ eigenvalues). Collecting all terms, we find the quantum capacity of the channel $\Lambda_{x}$ in closed form as
$Q(\Lambda)=\log d+\frac{d-1}{d}x\log\frac{x}{d}+(1-x+\frac{x}{d})\log(1-x+\frac{x}{d}).$ (21)
This is the exact capacity of the channel $\Lambda$ which smoothly drops from $\log_{2}d$ for the identity channel to 00 for the fully decohering channel, as expected. When $x=1$, the channel output is a classical mixture and has no superposition.
4 The Block-decohering channel $\Lambda_{k}$
We now extend this method to the channel in which decoherence occurs between subspaces of dimension $k$ and coherence in each subspace is maintained. Such a channel is defined by the map
$\Lambda_{k}(\rho)=(1-x)\rho+xD_{k}(\rho).$ (22)
Here $D_{k}$ depicts the decoherence map between blocks
$D_{k}(\rho)=\begin{pmatrix}\rho_{1}&\cdot&\cdot&\cdot&\cdot\\ \cdot&\rho_{2}&\cdot&\cdot&\cdot\\ \cdot&\cdot&\rho_{3}&\cdot&\cdot\\ \cdot&\cdot&\ &\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\rho_{r}\end{pmatrix}$ (23)
where each $\rho_{i}$ is the projector of the density matrix onto the $i-th$ subspace of dimension $k$. The dimension of the full Hilbert space is taken to be $d=kr$, where $k$ is the dimension of each subspace and $r$ is the number of orthogonal subspaces, i.e. number of blocks. As a map $D_{k}(\rho)$ is given by
$D_{k}(\rho)=\sum_{i=1}^{r}\pi_{i}\rho\pi_{i},$ (24)
where $\pi_{1}=|1\rangle\langle 1|+\cdots|k\rangle\langle k|$, $\pi_{2}=|k+1\rangle\langle k+1|+\cdots|2k\rangle\langle 2k|$ and so on. The complementary channel now becomes
$\Lambda_{k}^{c}(\rho)=\begin{pmatrix}(1-x)Tr(\rho)&aTr(\rho_{1})&aTr(\rho_{2})&\cdot&aTr(\rho_{r})\\ aTr(\rho_{1})&xTr(\rho_{1})&0&0&0\\ aTr(\rho_{2})&0&xTr(\rho_{2})&0&0\\ \cdot&0&0&\cdot&0&\\ aTr(\rho_{r})&0&0&0&xTr(\rho_{r}),\end{pmatrix}$ (25)
where $a=\sqrt{x(1-x)}$. Both of the basic properties (3) and (7) hold also for this channel from which we find that this channel is also degradable. Moreover, the Schur-Horn theorem is used to show that optimization can be performed by searching only among the diagonal density matrices $\rho_{d}$. The channel now has an extra symmetry, that is it is covariance under two types of unitary operations, that is:
$\Lambda_{k}(\mathbb{X}\rho\mathbb{X})=\mathbb{X}\Lambda_{k}(\rho)\mathbb{X},\hskip 28.45274pt\Lambda_{k}(\Omega\rho\Omega)=\Omega\Lambda_{k}(\rho)\Omega,$ (26)
where
$\mathbb{X}=\begin{pmatrix}X_{k}&\cdot&\cdot&\cdot&\cdot\\ \cdot&X_{k}&\cdot&\cdot&\cdot\\ \cdot&\cdot&X_{k}&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&X_{k}\end{pmatrix}$ (27)
which affects shift in each subspace and
$\Omega=\begin{pmatrix}\cdot&\cdot&\cdot&\cdot&I_{k}\\ \ I_{k}&\cdot&\cdot&\cdot&\cdot\\ \cdot&I_{k}&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&I_{k}&\cdot\end{pmatrix}$ (28)
which affects shifts among subspaces. We now use this symmetry to show that for an arbitrary diagonal matrix $\rho_{d}$,
$\frac{1}{d}\sum_{i=1}^{k}\sum_{j=1}^{r}U_{ij}\rho_{d}U_{ij}^{\dagger}=\frac{I}{d},$ (29)
where $U_{ij}=\mathbb{X}^{i}\ \Omega^{j}$. In the same way as for the channel $\Lambda_{1}$, concavity of the coherent information now yields that
$\max_{\rho}I_{c}(\Lambda_{k},\rho)=S(\Lambda_{k}(\frac{I}{d}))-S(\Lambda_{k}^{c}(\frac{I}{d}))$ (30)
where $d=rk$ is the dimension and $\Lambda_{k}^{c}$ has exactly the same form as (21), with $d$ replaced by $r$. The final result is given by
$Q(\Lambda_{k})=\log k+\log r+\frac{r-1}{r}x\log\frac{x}{r}+(1-x+\frac{x}{r})\log(1-x+\frac{x}{r}).$ (31)
The capacity of the channels at $x=0$ is given by $Q(\Lambda(x=0))=\log_{2}d$ and drops to $Q(\Lambda(x=1))=\log_{2}k$ This is a reflection of the fact that we can encode the input states into those subspaces which are not decohered. Plots of capacities of these channels is shown in figure (1) for the case where dimension $d=12$ and the number of blocks increase from $k=1$ (complete coherence) to $k=12$ (no coherence).
Figure 1: The quantum capacities of the block-decohering channels $\Lambda_{k}$ for $d=12$ for various values of $k$. For each $k$, the number of blocks is given by $r=\frac{12}{k}$. It is seen that even when $x=1$, full decoherence among subspaces, allows a nonzero quantum capacity. This is the capacity of quantum information transfer, when a suitable encoding into decoherence-free subspaces is used.
5 The Weakly-Decohering Channel $\Phi_{k}$
The non-zero capacity of block-decohering channels in the previous section may be attributed to the presence of decoherence-free subspaces, namely the $k-$ dimensional decoherence-free subspaces which are not affected by the decohering noise. As we will see in this section, in the absence of these decoherence-free subspaces, and provided that some degree of coherence exists, the channel can still have non-zero quantum capacity. In this way, we are considering a noise which is not so strong to annihilate all the off-diagonal elements. Rather, we assume an equal number of minor diagonals above and below the main diagonal are also kept intact while all the other off-diagonal elements vanish. As an example, for $d=6$, an example of such weak decoherence is the following:
$\rho\longrightarrow W_{2}(\rho):=\begin{pmatrix}\rho_{11}&\rho_{12}&0&0&0&\rho_{16}\\ \rho_{21}&\rho_{22}&\rho_{23}&0&0&0\\ 0&\rho_{32}&\rho_{33}&\rho_{34}&0&0\\ 0&0&\rho_{43}&\rho_{44}&\rho_{45}&0\\ 0&0&0&\rho_{54}&\rho_{55}&\rho_{56}\\ \rho_{61}&0&0&0&\rho_{65}&\rho_{66}\end{pmatrix}.$
Note that for simplicity of calculation we assume a periodic boundary condition on the states, i.e. $|d+1\rangle\equiv|1\rangle$, that is why in this example the two elements in the corner are non-vanishing. There is a related concept of coherence, called $k-$ coherence [36, 25], which is somewhat more general than this, but we will not consider it here. The original notion of incoherence [58], which defines incoherent states as diagonal density matrices in a specific basis, has been generalized to multi-level or k-coherence [25]. In this generalized setting, a state $|\psi\rangle=\sum_{i=1}^{d}\psi_{i}|i\rangle$ is said to have coherence at level k, if exactly $k$ of the coefficients $\psi_{i}$ are non-zero. Thus an incoherent state has coherence at level $1$, and a state like $|\psi\rangle=a|0\rangle+b|1\rangle$ in a $d-$ dimensional Hilbert space has coherence at level $2$ and so on. The generalization to mixed states is done by defining the states with coherence level $k$ to be the convex combination of all pure states whose coherence level is less than or equal to $k$. The notion of weak-coherence that we adopt here is different in that it assigns an order to the basis states. For example in a $d-$ dimensional space, if for some physical reason, say the photon number in each state, we order the states as $|0\rangle,|1\rangle,|2\rangle,\cdots$, two states of the form $|0\rangle+|1\rangle$ and $|1\rangle+|2\rangle$ are equally affected by decoherence, but a state like $|1\rangle+|3\rangle$ completely loses its coherence.
In this section, we consider decohering channels which with a certain probability decohere the input state to level-$k$ coherent states of the type we defined above. That is, we consider channels of the form
$\Phi_{k}(\rho)=(1-x)\rho+xW_{k}(\rho),$ (32)
where $W_{k}(\rho)$ is a weakly coherent density matrix at level $k$. Physically this means that the channel mitigates all coherences among the basis states which are further apart than $k$ in an ordered set of indices $\{|1\rangle,|2\rangle,\cdots|d\rangle\}$. The Kraus operators of this channel are:
$A_{0}=\sqrt{1-x}I,\hskip 28.45274ptA_{i}=\sqrt{\frac{x}{k}}\sum_{l=0}^{k-1}|i+l\rangle\langle i+l|.$ (33)
In other words, each $A_{i}$ is proportional to the projection operator on a $k-$ dimensional subspace, but this time, in contrast to the previous section, the subspaces are not orthogonal and they overlap on smaller subspaces. We first note that the basic property (3) still holds here, that is
$D\circ\Phi_{k}=\Phi_{k}\circ D=D.$ (34)
For the second property (7), we show that for the channel $\Phi_{k}$, the complementary channel $\Phi_{k}^{c}$ still depends on the diagonal elements of $\rho$. To see this, it is sufficient to note that the Kraus operators $I$, $A_{i}$ and the products $A_{i}A_{j}$ are all projectors on subspaces and hence extract the diagonal elements of $\rho$, when inserted in equation (5) for deriving the complementary channel. Therefore we have
$\Phi_{k}^{c}=\Phi_{k}^{c}\circ D.$ (35)
Combining these two equations, as before, we find
$\Phi_{k}^{c}\circ\Phi_{k}=(\Phi_{k}^{c}\circ D)\circ\Phi_{k}=\Phi_{k}^{c}\circ(D\circ\Phi_{k})=\Phi_{k}^{c}\circ D=\Phi_{k}$ (36)
which shows that the channel $\Phi_{k}$ is also degradable. Invoking Schur-Horn theorem as before, we need only search among the diagonal density matrices. We also note from (33) that $XA_{0}X^{\dagger}=A_{0},$ and $XA_{i}X^{\dagger}=A_{i+1}$, which show that the channel is covariant under shift operator $X$, that is
$\Phi_{k}(X\rho X^{\dagger})=X\Phi_{k}(\rho)X^{\dagger}.$ (37)
Combining these, we find from the concave property of the coherent information that the maximum value is obtained at $\rho_{d}=\frac{I}{d}$. Thus we have
$Q(\Phi_{k})=\max_{\rho_{d}}=I_{c}(\Phi_{k},\frac{I}{d}).$
To proceed we define
${\cal M}:=\Phi_{k}(\frac{I}{d})$
and calculate the following elements:
[…]
where due to the period condition of indices $s\equiv s+d$, we find after a close examination that $b_{s}$ is given by
$b(s):=\begin{cases}k-s&\text{if }\ \ 0\leq s\leq k-1,\\ 0&\text{if }\ \ k\leq s\leq d-k,\\ s+k-d&\text{if}\ \ d-k+1\leq s\leq d-1.\end{cases}$ (41)
Thereofore we find that
${\cal M}=\left[\begin{array}[]{c|c}1-x&\xi\sqrt{k}\bra{v_{0}}\\ \hline\cr\xi\sqrt{k}\ket{v_{0}}&xB\end{array}\right]$ (42)
where again $\xi=\sqrt{\frac{x(1-x)}{d}}$ and |$\ket{v_{0}}=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\ket{i}$. and the matrix $B$ is a circulant matrix given by
$B=\mathrm{Circ}(b_{0},b_{1},\dots,b_{d-1})=\begin{bmatrix}b_{0}&b_{1}&b_{2}&\cdots&b_{d-2}&b_{d-1}\\ \\ b_{d-1}&b_{0}&b_{1}&\cdots&b_{d-3}&b_{d-2}\\ \\ b_{d-2}&b_{d-1}&b_{0}&\cdots&b_{d-4}&b_{d-3}\\ \\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ \\ b_{1}&b_{2}&b_{3}&\cdots&b_{d-1}&b_{0}\end{bmatrix}.$ (43)
The matrix $B$, being circulant, has the structure
$B=\sum_{s=0}^{d-1}b_{s}X^{s},$
where $X$ is the shift operator. The eigenvalues of $X$ are roots of unity in the form $\omega^{m}$ corresponding to the eigenvector $|v_{m}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}\omega^{-im}|i\rangle$, ($m=0,1,\cdots d-1$), where $\omega=e^{\frac{2\pi i}{d}}$. Thus the eigenvalues of $B$ are of the form
$\lambda_{m}=\sum_{s=0}^{d-1}b_{s}\omega^{ms},\hskip 28.45274ptm=0,1,2,\cdots d-1.$ (44)
In view of (41), this turns out to be
$\lambda_{m}=\frac{1}{dk}\left[\frac{\sin{\left(\frac{k\pi m}{d}\right)}}{\sin{\left(\frac{\pi m}{d}\right)}}\right]^{2}.$ (45)
To find the quantum capacity of $\Phi_{k}$, we now need to diagonalize the matrix ${\cal M}\equiv\Phi_{k}^{c}(\frac{I}{d})$, as given in (42). We note that $|v_{0}\rangle$ is an eigenvector of $X$ and hence of $B$ with eigenvalue equal to $\lambda_{0}=\frac{k}{d}$. To find the spectrum of this matrix, we note that for any eigenvector $|v_{m\neq 0}\rangle$, the vector
$\begin{bmatrix}0\\ |v_{m}\rangle\end{bmatrix}$ (46)
is an eigenvector of ${\cal M}$ with eigenvalue equal to $x\lambda_{m}$. The other two eigenvectors lie in the span of |$\ket{v_{0}}$ and |$\ket{0}$:
${\cal M}(\alpha\ket{0}+\ket{v_{0}})=(\alpha(1-x)+d\zeta)\ket{0}+(\alpha\zeta+\frac{k}{d}x)\ket{v_{0}}$ (47)
The corresponding eigenvalue equation becomes, the solution of which yields two eigenvalues, namely
$\{0,\ \ \frac{kx}{d}+(1-x)\}.$
As a result, the set of eigenvalues of ${\cal M}=\Phi_{k}^{c}(\frac{I}{d})$ is:
$\{x\lambda_{1},x\lambda_{2},\ldots,x\lambda_{d-1}\}\cup\left\{0,\frac{kx}{d}+(1-x)\right\}$
where $\lambda_{m}$ is given in (45). This leads to the final expression of the quantum capacity, namely We can now evaluate the quantum capacity of $\Phi$:
$Q(\Phi_{k})=I_{c}\left(\Phi_{k},\frac{I}{d}\right)=\log{d}+\sum_{i=1}^{d-1}x\lambda_{i}\log(x\lambda_{i})+\left[\frac{kx}{d}+(1-x)\right]\log{\left[\frac{kx}{d}+(1-x)\right]}$ )
Figure 2: Coherent information $\lambda\bigl{(}\Phi,\tfrac{I}{d}\bigr{)}$ versus $x$ for $d=12$ and $k\in\{1,2,3,4,6\}$.
As special cases we note that when $k=1$, this capacity reduces to that of (21), as expected. Also for $x=0$, or when $k=d$, for which the channel is the identity channel, the capacity is given by $Q(\Phi_{k})=\log_{2}d.$ Note that for the strongest level of noise $(x=1)$, the quantum capacity is no longer given by $\log_{2}k$. The capacities of these channels for $d=12$ and for various degrees of decoherence, determined by $k$ are plotted in figure (2). It is worth nothing that even at full decoherence, the channel has non-zero capacity, despite having no decoherence-free subspace.
6 Discussion
In this work, we have derived an exact, closed-form expression for the quantum capacity of a wide class of decohering channels in arbitrary dimensions. These channels, which interpolate between ideal transmission and complete decoherence, are shown to be degradable for all noise parameters—allowing their capacity to be computed without regularization. This result significantly expands the set of known channels with exactly computable capacities, particularly in higher dimensions where such cases are extremely rare.
Beyond the exact formulas, our findings reveal a deeper structural insight: the geometry of decoherence—specifically the structure of coherence-preserving subspaces—plays a decisive role in determining whether a channel can sustain quantum information transmission, see also [59]. We show that even fully decohering channels may retain nonzero quantum capacity, provided that the decoherence acts in a way that respects overlap between certain subspaces. This challenges the conventional assumption that maximal decoherence necessarily eliminates all capacity.
The channels we analyze may serve as analytically tractable models for noise processes in a variety of physical platforms, including quantum memories, spin chains, and systems with partial access to coherence-preserving degrees of freedom. Our results therefore provide a bridge between the theoretical structure of quantum capacity and experimentally relevant noise models, with implications for designing robust encoding strategies in the presence of structured decoherence. These exact solutions open avenues for designing noise-resilient quantum memories and networks, bridging fundamental limits with engineering constraints in the NISQ era [60, 61, 62, 63]. Finally, our work opens up several new directions in the problem of capacity of quantum channels. The first one is to use these results for obtaining upper or lower bounds for more general models of noise in arbitrary dimensions. The second avenue is based on extending our method for proving degradability to more general channels.
Acknowledgement: We thank Stefano Chessa for his careful reading of the manuscript and for helpful correspondence, including drawing our attention to some relevant works. V.K. Thanks Dariusz Chruscinski, Saverio Pascazio, Gniewko Sarbicki and Mario Ziman for their valuable comments, when this work was presented in the 56th Symposium in Mathematical Physics in Torun, Poland.